Suppose you toss a six-sided die 10 times. What is the probability of never rolling a 4?
Answer #1

Let Y denotes the number of events of getting 4 in the roll of six-sided die 10 times. The random variable Y follows binomial distribution with parameters N = 10 and p = 1 / 6.

The probability distribution of Y is given as small Pleft ( Y=y ight )=10C_{y}left ( 1/6 ight )^{y}(5/6)^{10-y} for y = 0, 1, ... 10.

Probability of never rolling a 4 in tossing six-sided die 10 times is given as

P (Y = 0) = 1000 (1/6)0 (5/6)10-0 = (5/6),-0.162

Probability of never rolling a 4 in tossing six-sided die 10 times is 0.162.